# Safe Navigation in Uncertain Crowded Environments Using Risk Adaptive CVaR Barrier Functions

### v4

@@ I Introduction
Safe navigation in crowded environments with dynamic obstacles remains a fundamental challenge in robotics due to obstacle uncertainty. The common approaches involve using risk metrics to quantify and enforce safety constraints under this uncertainty, such as conventional stochastic control methods [1, 2] and robust control methods [3, 4]. While robust control methods prioritize worst-case scenarios, often leading to overly conservative behaviors, stochastic methods optimize for expected performance, which may result in unsafe decisions in high-risk situations.
Recently, [5, 4, 6] have explored combining Conditional Value-at-Risk (CVaR) with Control Barrier Functions (CBFs) to achieve probabilistic safety, thereby providing a trade-off between conservatism and efficiency. In their frameworks, CBFs [7] act as safety filters that regulate control inputs, while CVaR [8] quantifies the expected risk under a given risk level. However, existing methods based on the CVaR barrier function (CVaR-BF) primarily rely on a fixed risk level that remains constant throughout the trajectory. This can limit their applicability in dynamic environments with crowded obstacles. We highlight that such a fixed risk level is not flexible enough: A low risk tolerance enhances safety but can render the optimization infeasible, whereas a high risk tolerance improves feasibility at the expense of safety.
To address this limitation, we propose a risk-adaptive navigation approach, a novel extension of the CVaR-BF framework that dynamically adjusts the risk level to maintain safety while ensuring trajectory feasibility. Instead of relying on a pre-specified risk parameter, our adaptive controller initializes the system with a conservative risk level (e.g., zero) and incrementally increases it until an optimization-feasible solution is achieved. In highly dynamic scenarios, where obstacles move unpredictably and rapidly, the robot requires sufficient time and space to respond and adjust its risk level. However, overly conservative strategies can limit feasible solutions in crowded environments [9, 10], so an approach that maintains safety without excessively restricting the decision space is essential. To this end, we introduce the concept of a “Dynamic Zone”, where the original safety distance is adaptively expanded based on the relative position and velocity between the robot and its surrounding obstacles. The robot adjusts its trajectory before nearing obstacles, but only when needed to avoid unnecessary conservatism, while also extending the risk-adaptive range.
@@ II-1 Risk-Aware Control under Uncertainties
Risk-aware control under uncertainties can be approached using a variety of risk measures. The approach in [1] achieves finite-time risk bounds by considering expected risk and penalizing collisions in the cost function. Similarly, [2, 11] develop probabilistic safety bounds over a finite time horizon using a discrete-time CBF condition. However, these approaches prioritize average behavior and do not provide safety guarantee at each time step. Recently, distributionally robust optimization has been designed to enhance the safety by considering a set of possible distributions to optimize decisions for worst-case scenarios. [4] reformulates the collision avoidance problem by computing safe halfspaces from obstacle sample trajectories via distributionally-robust optimization. [3] considers the reachable set of the states of an obstacle under worst-case noise. Although these methods offer robust safety guarantees, they tend to be overly conservative, which can lead to infeasible solutions in dense environments.
@@ II-2 Dynamic Obstacle Avoidance
For safety-critical navigation, a rigorous theoretical framework and analysis are required. In [15], a coupled planning approache is proposed to model the obstacles’ motion and solve a joint optimization problem with explicit safety constraints. However, coupled methods falter when the obstacle model is inaccurate [16], and accurately modeling obstacle motion is challenging. CBFs [7] have been developed as safety filters that mitigate unsafe control inputs. Some approaches partition the solution space into convex regions using separating hyperplanes [17, 4]. For example, [4] reformulates the collision avoidance problem by computing safe halfspaces based on dynamic obstacle sample trajectories. Recent work has proposed a CBF approach combined with velocity obstacles [9, 10], where the velocity obstacle principle allows the vehicle to maintain forward motion without compromising safety. However, this method tends to trigger avoidance maneuvers even when obstacle are far away, thereby leading to an conservative behavior.
@@ III-B Probabilistic Constraints
Due to the inherent uncertainty in the dynamic obstacle scenarios, the safe states of the robot cannot be computed deterministically. Therefore, a probabilistic formulation of safety is required. The estimated obstacle state at time $k$, denoted $\hat{\mathbf{x}}^{o}_{k}\in\mathcal{O}\subset\mathbb{R}^{n}$, is obtained via sensor data and state estimation, and is treated as a deterministic value in our framework. However, $\mathbf{x}^{o}_{k+1}$ is modeled as a random variable, reflecting the inherent uncertainty in the predicted motion of the obstacle. Define the safe set as
@@
It is shown that (8) can be solved by the following optimization problem
@@ III-C Discrete-Time CVaR Barrier Functions
To handle uncertainty in a risk-aware manner, we employ a dynamic coherent risk measure known as CVaR-Safety [5]. Define the cumulative CVaR values from time 00 to $k$ as the composition of per-step CVaR operators applied sequentially over the time horizon. Formally, it is defined as:
@@
For the discrete-time system (1) and a risk level $\beta\in(0,1)$, a continuous function $h:\mathbb{R}^{n}\to\mathbb{R}$ is called a CVaR barrier function for the safe set $\mathcal{S}$ (5) if there exists a constant $\gamma\in(0,1]$ such that for each $\mathbf{x}_{k}\in\mathbb{R}^{n}$, there exist a $\mathbf{u}_{k}\in\mathbb{R}^{m}$ such that:
@@
Proof.
The proof can be found in [5]. ∎
Figure 2: Illustration of relationship between CVaR, the risk level $\beta$, and safety. The shaded blue region represents the boundary of $\operatorname{CVaR}_{\beta}(h_{k+1})$, which includes: (a) all possible obstacle positions, (b) the portion of the distribution captured by the CVaR at the specified risk level $\beta$, and (c) the expected (mean) positions of the obstacles. Here, $p$ refers to a Probabilistic Mass Function (PMF).
@@ IV-A Safety and Feasibility Analysis
Subsequently, at a given risk level $\beta$, the set of admissible control inputs that satisfy the CVaR-BF constraint at time $k$ is defined as
@@ IV-B Adaptive Risk Level
Consider the discrete-time system (1) and an adaptive risk level $\beta_{k}$ at each time step $k$ as defined in (18). A function $h:\mathbb{R}^{n}\to\mathbb{R}$ is called a Risk Adaptive CVaR Barrier Function for the safe set $\mathcal{S}$ in (5) if there exists a constant $\gamma\in(0,1]$ such that for each $\mathbf{x}_{k}\in\mathbb{R}^{n}$, there exist a $\mathbf{u}_{k}\in\mathbb{R}^{m}$ such that,
@@ V Dynamic Zones for CVaR Barrier Functions
Effective risk management is essential for robotic navigation, particularly in environments with high-speed obstacles and uncertainties that shorten reaction times and increase collision risks. Fig. 1a shows that using a fixed risk level can render the problem infeasible when the robot approaches an obstacle. In contrast, Fig. 1b employs an adaptive risk level (as described in Sec. IV-B) that starts conservatively prompting the robot to initiate a turn early, and then relaxes the safety requirements as needed. However, without incorporating a dynamic zone that provides a virtual radius (an extra buffer), this risk adjustment can ultimately compromise safety and lead to collisions. Therefore, as shown in Fig. 1c, the combination of an adaptive risk level with a dynamic zone not only facilitates early obstacle avoidance but also provides greater flexibility for risk adjustments, thereby relaxing constraints and expanding the feasible space.
@@ V-A Dynamic Zone–Based Barrier Function
where $\phi$ is the half-angle of the collision cone, defined as $\cos\phi=\sqrt{\|\mathbf{p}^{\text{rel}}_{k}\|^{2}-R_{\text{safe}}^{2}}\,/\,\|\mathbf{p}^{\text{rel}}_{k}\|$, and $\mathbf{v}_{k}\in\mathbb{R}^{d}$ and ${\color[rgb]{0,0,0}\hat{\mathbf{v}}^{o}_{k}}\in\mathbb{R}^{d}$ represent the velocity of robot and obstacle, respectively. This CBF enforces that the angle between $\mathbf{p}^{\text{rel}}_{k}$ and $\mathbf{v}^{\text{rel}}_{k}$ remains less than $\pi-\phi$, thereby ensuring that the robot is directed away from the obstacle. However, such geometric constraints can be overly conservative. In highly dynamic environments, they may force the robot to execute unnecessarily, leading to premature path diversions and, in some cases, rendering the navigation problem infeasible [10].
@@
  • If $\langle\mathbf{p}^{\text{rel}}_{k},\mathbf{v}^{\text{rel}}_{k}\rangle\geq 0$Rightarrow $\Delta_{k}=0$. This implies that the robot and the obstacle are moving away from each other, thereby reducing the likelihood of a collision.
  • If $\langle\mathbf{p}^{\text{rel}}_{k},\mathbf{v}^{\text{rel}}_{k}\rangle<0$, then $\Delta_{k}>0$, indicating that the robot and the obstacle are approaching each other, thereby increasing the likelihood of a collision.
@@ V-B Probabilistic Safety Guarantee
Next, we detail how to derive a new upper bound, denoted by $\bar{\beta}_{u}$, for the adaptive risk level within the dynamic zone–based approach. The key insight is that the dynamic zone represents a larger, dynamic, yet virtual safety distance, rather than the actual physical separation between the robot and an obstacle (see Fig. 1c). We propose an analytical formulation that leverages this expanded safety zone to permit a wider range of risk level adaptation, i.e., $\beta_{k}\in(0,\bar{\beta}_{u}]$, while preserving the same probabilistic safety guarantee as that achieved with the original conservative risk level $\beta_{u}$ associated with the conventional distance-based barrier function $h^{\text{D}}_{k}$ (20).
Let $h^{\text{D}}_{k+1}$ denote the original distance‐based function with probability density function $P_{h^{\text{D}}_{k+1}}(x)$, where $x$ represents a possible realization of $h^{\text{D}}_{k+1}$. Recall that our dynamic zone‐based barrier function $h^{\text{Z}}_{k+1}$ is defined by expanding the safety radius via a factor related to $\Delta_{k}$; equivalently, we express it as a deterministic shift:
@@
Since this shift is deterministic, it does not change the shape of the distribution of $h^{\text{D}}_{k+1}$. Thus, the probability density of $h^{\text{Z}}_{k+1}$ satisfies
$P_{h^{\text{Z}}_{k+1}}(x)=P_{h^{\text{D}}_{k+1}}\Bigl{(}x+\Delta_{k+1}\,R_{\text{safe}}^{2}\Bigr{)}.$ (24)
Moreover, by the translation invariance property of CVaR [22], we have
$\operatorname{CVaR}^{k}_{\beta}\bigl{(}h^{\text{Z}}_{k+1}\bigr{)}=\operatorname{CVaR}^{k}_{\beta}\bigl{(}h^{\text{D}}_{k+1}\bigr{)}-\Delta_{k+1}\,R_{\text{safe}}^{2}.$ (25)
@@
$\operatorname{CVaR}^{k}_{\beta}\bigl{(}h^{\text{Z}}_{k+1}\bigr{)}\geq(1-\gamma)\,h^{\text{Z}}_{k}\Leftrightarrow$ (26)
$\operatorname{CVaR}^{k}_{\beta}\bigl{(}h^{\text{D}}_{k+1}\bigr{)}\geq(1-\gamma)h^{\text{D}}_{k}+\bigl{(}\Delta_{k+1}-(1-\gamma)\Delta_{k}\bigr{)}R_{\text{safe}}^{2}.$
Since CVaR is monotonic with respect to the risk level, this additional term allows us to define a new, less conservative upper bound on the risk level. One can show that
$\operatorname{CVaR}^{k}_{\bar{\beta}_{u}}\bigl{(}h^{\text{Z}}_{k+1}\bigr{)}=$ $\operatorname{CVaR}^{k}_{\beta_{u}}\bigl{(}h^{\text{D}}_{k+1}\bigr{)}+$ (27)
$\bigl{(}\Delta_{k+1}-(1-\gamma)\Delta_{k}\bigr{)}R_{\text{safe}}^{2},$
where $\beta_{u}$ is the original risk level upper bound. Thus, by adopting the dynamic zone-based barrier function $h^{\text{Z}}_{k+1}$ and selecting the adaptive risk level according to
@@
Under the distance‐based formulation, the adaptive risk level $\beta_{k}$ is chosen from $(0,\beta_{u}]$ so that the risk adaptive CVaR-BF guarantees safety (by Theorem (2)). In the dynamic zone–based formulation, $\beta_{k}$ is selected from $(0,\bar{\beta}_{u}]$. By design, when the adaptive risk level in the dynamic zone approach reaches its upper bound $\bar{\beta}_{u}$, the resulting CVaR condition is equivalent to that obtained with the upper bound $\beta_{u}$ in the distance‐based case. Therefore, the safe set $\mathcal{S}$ remains CVaR‐safe with at least the same probabilistic safety guarantee. ∎

### v5

@@ I Introduction
Safe navigation in crowded environments with dynamic obstacles remains a fundamental challenge in robotics due to obstacle uncertainty. The common approaches involve using risk metrics to quantify and enforce safety constraints under this uncertainty, such as conventional stochastic control methods [1, 2, 3] and robust control methods [4, 5]. While robust control methods prioritize worst-case scenarios, often leading to overly conservative behaviors, stochastic methods optimize for expected performance, which may result in unsafe decisions in high-risk situations.
The combination of Conditional Value-at-Risk (CVaR) and Control Barrier Functions (CBFs) has recently been explored to achieve probabilistic safety, offering a trade-off between conservatism and efficiency [6, 5, 7]. In these frameworks, CBFs [8, 9] act as safety filters that regulate control inputs, while CVaR [10] quantifies the expected risk under a given risk level. However, existing methods based on the CVaR barrier function (CVaR-BF) primarily rely on a fixed risk level that remains constant throughout the trajectory. This can limit their applicability in dynamic environments with crowded obstacles. We highlight that such a fixed risk level is not flexible enough: A low risk tolerance enhances safety but can render the optimization infeasible, whereas a high risk tolerance improves feasibility at the expense of safety.
To address this limitation, we propose a risk-adaptive navigation approach, a novel extension of the CVaR-BF framework that dynamically adjusts the risk level to maintain safety while ensuring trajectory feasibility. Instead of relying on a pre-specified risk parameter, our adaptive controller initializes the system with a conservative risk level (e.g., zero) and incrementally increases it until an optimization-feasible solution is achieved. In highly dynamic scenarios, where obstacles move unpredictably and rapidly, the robot requires sufficient time and space to respond and adjust its risk level. However, overly conservative strategies can limit feasible solutions in crowded environments [11, 12], so an approach that maintains safety without excessively restricting the decision space is essential. To this end, we introduce the concept of a “Dynamic Zone", where the original safety distance is adaptively expanded based on the relative position and velocity between the robot and its surrounding obstacles. The robot adjusts its trajectory before nearing obstacles, but only when needed to avoid unnecessary conservatism, while also extending the risk-adaptive range.
@@ II-1 Risk-Aware Control under Uncertainties
Risk-aware control under uncertainties can be approached using a variety of risk measures. The approach in [1] achieves finite-time risk bounds by considering expected risk and penalizing collisions in the cost function. Similarly, [2, 13] develop probabilistic safety bounds over a finite time horizon using a discrete-time CBF condition. However, these approaches prioritize average behavior and do not provide safety guarantee at each time step. Recently, distributionally robust optimization has been designed to enhance the safety by considering a set of possible distributions to optimize decisions for worst-case scenarios. [5] reformulates the collision avoidance problem by computing safe half-spaces from obstacle sample trajectories via distributionally-robust optimization. [4] considers the reachable set of the states of an obstacle under worst-case noise. Although these methods offer robust safety guarantees, they tend to be overly conservative, which can lead to infeasible solutions in dense environments.
@@ II-2 Dynamic Obstacle Avoidance
For safety-critical navigation, a rigorous theoretical framework and analysis are required. In [17], a coupled planning approach is proposed to model the obstacles’ motion and solve a joint optimization problem with explicit safety constraints. However, coupled methods falter when the obstacle model is inaccurate [18], and accurately modeling obstacle motion is challenging. CBFs [8, 9] have been developed as safety filters that mitigate unsafe control inputs. Some approaches partition the solution space into convex regions using separating hyperplanes [19, 5]. For example, [5] reformulates the collision avoidance problem by computing safe half-spaces based on dynamic obstacle sample trajectories. Recent work has proposed a CBF approach combined with velocity obstacles [11, 12, 20], where the velocity obstacle principle allows the vehicle to maintain forward motion without compromising safety. However, this method tends to trigger avoidance maneuvers even when obstacle are far away, thereby leading to an conservative behavior.
@@ III-B Probabilistic Constraints
Due to the inherent uncertainty in the dynamic obstacle scenarios, the safe states of the robot cannot be computed deterministically. Therefore, a probabilistic formulation of safety is required. The estimated obstacle state at time $k$, denoted $\hat{\mathbf{x}}^{o}_{k}\in\mathcal{O}\subset\mathbb{R}^{n}$, is obtained via sensor data and state estimation, and is treated as a deterministic value in our framework. However, the future state of the obstacle $\mathbf{x}^{o}_{k+1}$ is modeled as a random variable, reflecting the inherent uncertainty in the predicted motion of the obstacle. Define the safe set as
@@
As demonstrated in [22], the expression in (8) can be reformulated as the following optimization problem
@@ III-C Discrete-Time CVaR Barrier Functions
To handle uncertainty in a risk-aware manner, we employ a dynamic coherent risk measure known as CVaR-Safety [6]. Define the cumulative CVaR values from time 0 to $k$ as the composition of per-step CVaR operators applied sequentially over the time horizon. Formally, it is defined as:
@@
For the discrete-time system (1) and a risk level $\beta\in(0,1)$, a continuous function $h:\mathbb{R}^{n}\to\mathbb{R}$ is called a CVaR barrier function for the safe set $\mathcal{S}$ (5) if there exists a constant $\gamma\in(0,1]$ such that for each $\mathbf{x}_{k}\in\mathbb{R}^{n}$, there exist a $\mathbf{u}_{k}\in\mathbb{R}^{m}$ s.t.:
@@
Figure 2: Illustration of relationship between CVaR, the risk level $\beta$, and safety. The shaded blue region represents the boundary of $\operatorname{CVaR}_{\beta}(h_{k+1})$, which includes: (a) all possible obstacle positions, (b) the portion of the distribution captured by the CVaR at a specified risk level $\beta$, and (c) the expected (mean) positions of the obstacles. Here, $p(h_{k+1})$ denotes the probability distribution of the $h_{k+1}$.
@@ IV-A Safety and Feasibility Analysis
At a given risk level $\beta$, the set of admissible control inputs that satisfy the CVaR-BF constraint at time $k$ is defined as
@@ IV-B Adaptive Risk Level
Consider the discrete-time system (1) and an adaptive risk level $\beta_{k}$ at each time step $k$ as defined in (18). A function $h:\mathbb{R}^{n}\to\mathbb{R}$ is called a Risk Adaptive CVaR Barrier Function for the safe set $\mathcal{S}$ in (5) if there exists a constant $\gamma\in(0,1]$ such that for each $\mathbf{x}_{k}\in\mathbb{R}^{n}$, there exist a $\mathbf{u}_{k}\in\mathbb{R}^{m}$ s.t.,
@@ V Dynamic Zones for CVaR Barrier Functions
Effective risk management is essential for robotic navigation, particularly in environments with high-speed obstacles and uncertainties that shorten reaction times and increase collision risks. Figure 1a shows that using a fixed risk level can render the problem infeasible when the robot approaches an obstacle. In contrast, Fig. 1b employs an adaptive risk level (as described in Sec. IV-B) that starts conservatively prompting the robot to initiate a turn early, and then relaxes the safety requirements as needed. However, without incorporating a dynamic zone that provides a virtual radius (an extra buffer), this risk adjustment can ultimately compromise safety and lead to collisions. Therefore, as shown in Fig. 1c, the combination of an adaptive risk level with a dynamic zone not only facilitates early obstacle avoidance but also provides greater flexibility for risk adjustments, thereby relaxing constraints and expanding the feasible space.
@@ V-A Dynamic Zone–Based Barrier Function
where $\phi$ is the half-angle of the collision cone, defined as $\cos\phi=\sqrt{\|\mathbf{p}^{\text{rel}}_{k}\|^{2}-R_{\text{safe}}^{2}}\,/\,\|\mathbf{p}^{\text{rel}}_{k}\|$, and $\mathbf{v}_{k}\in\mathbb{R}^{d}$ and ${\color[rgb]{0,0,0}\hat{\mathbf{v}}^{o}_{k}}\in\mathbb{R}^{d}$ represent the velocity of robot and obstacle, respectively. This CBF enforces that the angle between $\mathbf{p}^{\text{rel}}_{k}$ and $\mathbf{v}^{\text{rel}}_{k}$ remains less than $\pi-\phi$, thereby ensuring that the robot is directed away from the obstacle. However, such geometric constraints can be overly conservative. In highly dynamic environments, they may force the robot to take unnecessary actions, leading to premature path diversions and, in some cases, rendering the navigation problem infeasible [12].
@@
  • If $\frac{\langle\mathbf{p}^{\text{rel}}_{k},\mathbf{v}^{\text{rel}}_{k}\rangle}{\|\mathbf{p}^{\text{rel}}_{k}\|\,\|\mathbf{v}^{\text{rel}}_{k}\|}\geq 0$Rightarrow $\Delta_{k}=0$. This implies that the robot and the obstacle are moving away from each other, thereby reducing the likelihood of a collision.
  • If $\frac{\langle\mathbf{p}^{\text{rel}}_{k},\mathbf{v}^{\text{rel}}_{k}\rangle}{\|\mathbf{p}^{\text{rel}}_{k}\|\,\|\mathbf{v}^{\text{rel}}_{k}\|}<0$Rightarrow $\Delta_{k}>0$. This implies that the robot and the obstacle are approaching each other, thereby increasing the likelihood of a collision.
@@ V-B Probabilistic Safety Guarantee
Next, we detail how to derive a new upper bound, denoted by $\bar{\beta}_{u}\in(0,1)$, for the adaptive risk level within the dynamic zone–based approach. The key insight is that the dynamic zone represents a larger, dynamic, yet virtual safety distance, rather than the actual physical separation between the robot and an obstacle (see Fig. 1c). We propose an analytical formulation that leverages this expanded safety zone to permit a wider range of risk level adaptation, i.e., $\beta_{k}\in(0,\bar{\beta}_{u}]$, while preserving the same probabilistic safety guarantee as that achieved with the original conservative risk level $\beta_{u}$ associated with the conventional distance-based barrier function $h^{\text{D}}_{k}$ (20).
Let $h^{\text{D}}_{k+1}$ denote the original distance-based function with probability density function $p_{h^{\text{D}}_{k+1}}(x)$, where $x$ represents a possible realization of $h^{\text{D}}_{k+1}$. Recall that our dynamic zone-based barrier function $h^{\text{Z}}_{k+1}$ is defined by expanding the safety radius via a factor related to $\Delta_{k+1}$; equivalently, we express it as a shift:
@@
By the subadditivity property of CVaR [24], we have
$\operatorname{CVaR}^{k}_{\beta}\bigl{(}h^{\text{Z}}_{k+1}\bigr{)}$ $=\operatorname{CVaR}^{k}_{\beta}\bigl{(}h^{\text{D}}_{k+1}-\Delta_{k+1}\,R_{\text{safe}}^{2}\bigr{)}$ (24)
$\leq\operatorname{CVaR}^{k}_{\beta}\bigl{(}h^{\text{D}}_{k+1}\bigr{)}-\operatorname{CVaR}^{k}_{\beta}\bigl{(}\Delta_{k+1}\,R_{\text{safe}}^{2}\bigr{)}.$
@@
$\operatorname{CVaR}^{k}_{\bar{\beta}_{u}}\bigl{(}h^{\text{Z}}_{k+1}\bigr{)}$ $\geq(1-\gamma)\,h^{\text{Z}}_{k}\Rightarrow$ (25)
$\operatorname{CVaR}^{k}_{\bar{\beta}_{u}}\bigl{(}h^{\text{D}}_{k+1}\bigr{)}$ $\geq(1-\gamma)h^{\text{D}}_{k}$
$+\bigl{(}\operatorname{CVaR}^{k}_{\beta}\bigl{(}\Delta_{k+1}\bigr{)}-(1-\gamma)\Delta_{k}\bigr{)}R_{\text{safe}}^{2}.$
Define a dynamic risk offset at time $k+1$ as $\delta_{k+1}\coloneqq\bigl{(}\operatorname{CVaR}^{k}_{\beta}\bigl{(}\Delta_{k+1}\bigr{)}-(1-\gamma)\,\Delta_{k}\,\bigr{)}R_{\text{safe}}^{2}$. When an obstacle moves toward the robot, causing the likelihood of a collision to increase, i.e., $\operatorname{CVaR}^{k}_{\beta}\bigl{(}\Delta_{k+1}\bigr{)}\geq\mathbb{E}(\Delta_{k+1})>(1-\gamma)\Delta_{k}$, this implies that $\delta_{k+1}>0$. Since CVaR is monotonic with respect to the risk level, this additional term allows us to define a new, less conservative upper bound on the risk level. One can show that, if the same safety threshold $R_{\text{safe}}$ is maintained, the risk level $\bar{\beta}_{u}$ permitted in the dynamic zone-based formulation can be greater than that in the original distance-based formulation $\beta_{u}$, i.e., $\bar{\beta}_{u}\geq\beta_{u}$. The maximum risk level $\bar{\beta}_{u}$ can be computed by searching for the value such that the following condition holds:
$\bar{\beta}_{u}=\{\beta\,|\operatorname{CVaR}^{k}_{\beta}(h_{k+1}^{D})=\operatorname{CVaR}^{k}_{\beta_{u}}(h_{k+1}^{D})+\delta_{k+1}\}$ (26)
Thus, by adopting the dynamic zone-based barrier function $h^{\text{Z}}_{k+1}$ and selecting the adaptive risk level according to
@@
Under the distance-based formulation, the adaptive risk level $\beta_{k}$ is chosen from $(0,\beta_{u}]$ so that the risk adaptive CVaR-BF guarantees safety (by Theorem (2)). In the dynamic zone–based formulation, $\beta_{k}$ is selected from $(0,\bar{\beta}_{u}]$. By design, when the adaptive risk level in the dynamic zone approach reaches its upper bound $\bar{\beta}_{u}$, the resulting CVaR condition is equivalent to that obtained with the upper bound $\beta_{u}$ in the distance-based case. Therefore, the safe set $\mathcal{S}$ remains CVaR-safe with at least the same probabilistic safety guarantee. ∎